# Reconstructing Physics-Informed Machine Learning for Traffic Flow Modeling: a Multi-Gradient Descent and Pareto Learning Approach

### v1

@@ 3.1.2 Dual cone gradient descent
6    $g_{t}^{c}=\frac{\nabla\mathcal{L}_{b}(\theta_{t})}{\|\nabla\mathcal{L}_{b}(\theta_{t})\|}+\frac{\nabla\mathcal{L}_{r}(\theta_{t})}{\|\nabla\mathcal{L}_{r}(\theta_{t})\|}$
7    $g_{t}^{dual}=\frac{\langle g_{t}^{c},\nabla\mathcal{L}(\theta_{t})\rangle}{\|g_{t}^{c}\|^{2}}g_{t}^{c}$
8    $\theta_{t}=\theta_{t-1}-\lambda g_{t}^{dual}$
@@
Algorithm 2 DCGD (Center) [Hwang and Lim [2024]]
1 Require: learning rate $\lambda$, max epoch $T$, initial point $\theta_{0}$, gradient threshold $\varepsilon$, conflict threshold $\alpha$
@@
6   if $\nabla\mathcal{L}(\theta_{t})\notin\mathbf{K}^{*}$ then
7     $g_{t}^{dual}=\frac{1}{2}\nabla\mathcal{L}(\theta_{t})_{\|(\nabla\mathcal{L}_{r}(\theta_{t}))^{\bot}}+\frac{1}{2}\nabla\mathcal{L}_{\|(\nabla\mathcal{L}_{b}(\theta_{t}))^{\bot}}$
8   else
9     $g_{t}^{dual}=\nabla\mathcal{L}(\theta_{t})$
10   end if
11    $\theta_{t}=\theta_{t-1}-\lambda g_{t}^{dual}$
12 end for
Algorithm 3 DCGD (Average) [Hwang and Lim [2024]]
1 Require: learning rate $\lambda$, max epoch $T$, initial point $\theta_{0}$, gradient threshold $\varepsilon$, conflict threshold $\alpha$
2 for $t=1$ to $T$ do
4     break
5   end if
6   if $\nabla\mathcal{L}(\theta_{t})\in\mathbf{K}^{*}$ then
7     $g_{t}^{dual}=\nabla\mathcal{L}(\theta_{t})$
8   else if $\nabla\mathcal{L}(\theta_{t})\notin\mathbf{K}^{*}\text{ and }\langle\nabla\mathcal{L}(\theta_{t}),\nabla\mathcal{L}_{r}(\theta_{t})\rangle<0$ then
9     $g_{t}^{dual}=\nabla\mathcal{L}(\theta_{t})_{\|(\nabla\mathcal{L}_{r}(\theta_{t}))^{\bot}}$
10   else if $\nabla\mathcal{L}(\theta_{t})\notin\mathbf{K}^{*}\text{ and }\langle\nabla\mathcal{L}(\theta_{t}),\nabla\mathcal{L}_{b}(\theta_{t})\rangle<0$ then
11     $g_{t}^{dual}=\nabla\mathcal{(}\theta_{t})_{\|(\nabla\mathcal{L}_{b}(\theta_{t})^{\bot}}$
@@ 7     $g_{t}^{dual}=\nabla\mathcal{L}(\theta_{t})$
In Algorithm 2, we first calculate a normalized sum of two directions $g^{c}_{t}$. $\angle\bigl{(}g^{c}_{t},\mathcal{L}_{b}(\theta_{t})\bigr{)}=\angle\bigl{(}g^{c}_{t},\mathcal{L}_{r}(\theta_{t})\bigr{)}=\frac{\phi_{t}}{2}$ 555 we use $\angle\bigl{(}\mathbf{u},\mathbf{v}\bigr{)}$ to represent the angle between two vectors, mathematically: $\angle\bigl{(}\mathbf{u},\mathbf{v}\bigr{)}:=\arccos\!\biggl{(}\frac{\mathbf{u}\cdot\mathbf{v}}{\|\mathbf{u}\|\;\|\mathbf{v}\|}\biggr{)}$ (16) , and then, we project the total loss $\mathcal{L}(\theta_{t})$ onto $g^{c}_{t}$ to get the final common descent direction $g^{dual}_{t}$.
@@
$\nabla\mathcal{L}(\theta_{t})=\nabla\mathcal{L}_{r}(\theta_{t})+\nabla\mathcal{L}_{b}(\theta_{t})$ (17)
@@
$\nabla\mathcal{L}(\theta_{t})\in\mathbf{K}^{*}$ (18)
we use $\nabla\mathcal{L}(\theta_{t})$ as the update. Otherwise, we first project $\nabla\mathcal{L}(\theta_{t})$ onto two boundary $\nabla\mathcal{L}(\theta_{t})_{\|(\nabla\mathcal{L}_{r}(\theta_{t}))^{\bot}}$ and $\nabla\mathcal{L}_{\|(\nabla\mathcal{L}_{b}(\theta_{t}))^{\bot}}$ of $\mathbf{G}_{t}$, respectively. Then, calculate the arithmetic average as the update:
$g^{dual}_{t}=\frac{1}{2}\nabla\mathcal{L}(\theta_{t})_{\|(\nabla\mathcal{L}_{r}(\theta_{t}))^{\bot}}+\frac{1}{2}\nabla\mathcal{L}_{\|(\nabla\mathcal{L}_{b}(\theta_{t}))^{\bot}}$ (19)
This guarantees that the chosen descent direction reduces both $\mathcal{L}_{r}(\theta_{t})$ and $\nabla\mathcal{L}_{b}(\theta_{t})$ simultaneously.
In Algorithm 4, at iteration $t$, the dual gradient $g_{t}^{\mathrm{dual}}$ is constructed based on whether the current gradient lies outside the dual cone $\mathbf{K}^{*}$ and whether it conflicts with bound gradients $\mathcal{L}_{b}(\theta_{t})$ and $\mathcal{L}_{r}(\theta_{t})$:
$g_{t}^{\mathrm{dual}}=\begin{cases}\nabla\mathcal{L}(\theta_{t}),&\nabla\mathcal{L}(\theta_{t})\in\mathbf{K}^{*},\\[4.30554pt] \displaystyle\nabla\mathcal{L}(\theta_{t})_{\|(\nabla\mathcal{L}_{r}(\theta_{t}))^{\bot}},&\nabla\mathcal{L}(\theta_{t})\notin\mathbf{K}^{*},\ \langle\nabla\mathcal{L}(\theta_{t}),\nabla\mathcal{L}_{r}(\theta_{t})\rangle<0,\\[8.61108pt] \displaystyle\nabla\mathcal{(}\theta_{t})_{\|(\nabla\mathcal{L}_{b}(\theta_{t})^{\bot}},&\nabla\mathcal{L}(\theta_{t})\notin\mathbf{K}^{*},\ \langle\nabla\mathcal{L}(\theta_{t}),\nabla\mathcal{L}_{b}(\theta_{t})\rangle<0.\end{cases}$ (20)
@@
Assume that both loss functions, $\mathcal{L}_{b}(\cdot)$ and $\mathcal{L}_{r}(\cdot)$, are differentiable and the total gradient $\nabla\mathcal{L}(\cdot)$ is $L$-Lipschitz continuous with $L>0$. If $g_{t}^{\text{dual}}$ satisfies the following two conditions:
@@
    $2\langle\nabla\mathcal{L}(\theta_{t}),g_{t}^{\text{dual}}\rangle-\|g_{t}^{\text{dual}}\|^{2}\geq 0$,
@@
    There exists $M>0$ such that $\|g_{t}^{\text{dual}}\|\geq M\|\nabla\mathcal{L}(\theta_{t})\|$,
@@
$\frac{1}{T+1}\sum_{t=0}^{T}\|\nabla\mathcal{L}(\theta_{t})\|^{2}\leq\frac{2\left(\mathcal{L}(\theta_{0})-\mathcal{L}(\theta^{*})\right)}{\lambda M(T+1)}.$ (21)
@@ 3.2 Potential limitation of linear scalarization
In most contemporary PIML studies, the primary and most effective method for calculating the final total loss is through linear scalarization, as shown in the form illustrated in Eq 2. As is aware by many previous studies, the specific values of the coefficients $\alpha$ and $\beta$ cause great difficulties in fine-tuning, and since the value of the coefficients will not change during the training process, it may lead to more serious problems:
@@
    Linear scalarization constructs a combined loss function as a linear weighted sum:
    $\mathcal{L}_{\alpha}(\theta)=\alpha\mathcal{L}_{1}(\theta)+(1-\alpha)\mathcal{L}_{2}(\theta),\quad\alpha\in[0,1]$
@@
    $\nabla\mathcal{L}_{\alpha}(\theta^{*})=(1-\alpha)\nabla\mathcal{L}_{2}(\theta^{*})$
@@
    $\mathcal{L}_{\beta}(\theta)=\beta\mathcal{L}_{1}(\theta)+(1-\beta)\mathcal{L}_{2}(\theta),\quad\beta\in[0,1],$
@@
    However, certain minimizers obtained by MGDA cannot be theoretically achieved by optimizing the scalarized total loss for any linear scalarization weights.
@@ 4.1 PIML model based on the macroscopic traffic flow model
The loss-function design for LWR-PINN is shown in Eq 27, where $N_{o}$ and $N_{a}$ denote the numbers of observation and auxiliary points, respectively. For each observation point
@@
After tuning the hyperparameters, the experimental results are displayed in Table 5, where the LWR-PINN model refers to the model shown in Figure 3, which adopted a loss function as: $Loss_{\theta,\omega}^{LWR}=100\cdot MSE_{o}+500\cdot MSE_{a}$, for all other models, they treat the training process as a bi-objective optimization problem, the two objectives are $MSE_{o}$ and $MSE_{a}$. According to Table 5, the TMGD-LWR-PINN model and the models based on DCGDs exhibit performance similar to that of the baseline model, the LWR-PINN model, with the exception of the DCGD-AVER-LWR-PINN model, which shows a significantly larger speed $\mathbb{L}^{2}$ error compared to the baseline. We believe this may be attributed to the following reason. In PIML studies that involve a macroscopic traffic flow model, Lei et al. [2025b] reveal that high-resolution data is needed to capture the true pattern of the physics dynamics described by the LWR model and ARZ model, which means that the $\Delta x$ and $\Delta t$ in those data should be as small as possible, $\Delta x$ represents the minimum distance between two traffic sensors, and $\Delta t$ represents the minim time interval between consecutive measurements. However, the general traffic sensor data might not be high-resolution data since general traffic sensor data usually has a very large $\Delta x$ and $\Delta t$. As a result, the physics part of the PIML model may always provide a gradient that could degrade the model’s performance, leading to a failure of the total PIML model.

### v3

@@ 8    $\boldsymbol{\theta}_{t}=\boldsymbol{\theta}_{t-1}-\lambda g_{t}^{dual}$
6   if $\nabla\mathcal{L}(\boldsymbol{\theta}_{t})\notin\mathbf{K}^{*}$ then
7     $g_{t}^{dual}=\frac{1}{2}\nabla\mathcal{L}(\boldsymbol{\theta}_{t})_{\|(\nabla\mathcal{L}_{p}(\boldsymbol{\theta}_{t}))^{\bot}}+\frac{1}{2}\nabla\mathcal{L}_{\|(\nabla\mathcal{L}_{d}(\boldsymbol{\theta}_{t}))^{\bot}}$
8   else
9     $g_{t}^{dual}=\nabla\mathcal{L}(\boldsymbol{\theta}_{t})$
10   end if
11    $\boldsymbol{\theta}_{t}=\boldsymbol{\theta}_{t-1}-\lambda g_{t}^{dual}$
@@
Algorithm 3 DCGD (Average) [Hwang and Lim [2024]]
1Require: learning rate $\lambda$, max epoch $T$, initial point $\theta_{0}$, gradient threshold $\varepsilon$, conflict threshold $\alpha$
@@
3   if $\pi-\alpha<\phi_{t}\text{ or }\|\nabla\mathcal{L}(\boldsymbol{\theta}_{t})\|<\varepsilon$ then
@@
6   if $\nabla\mathcal{L}(\boldsymbol{\theta}_{t})\in\mathbf{K}^{*}$ then
7     $g_{t}^{dual}=\nabla\mathcal{L}(\boldsymbol{\theta}_{t})$
8   else if $\nabla\mathcal{L}(\boldsymbol{\theta}_{t})\notin\mathbf{K}^{*}\text{ and }\langle\nabla\mathcal{L}(\boldsymbol{\theta}_{t}),\nabla\mathcal{L}_{p}(\boldsymbol{\theta}_{t})\rangle<0$ then
9     $g_{t}^{dual}=\nabla\mathcal{L}(\boldsymbol{\theta}_{t})_{\|(\nabla\mathcal{L}_{p}(\boldsymbol{\theta}_{t}))^{\bot}}$
10   else if $\nabla\mathcal{L}(\boldsymbol{\theta}_{t})\notin\mathbf{K}^{*}\text{ and }\langle\nabla\mathcal{L}(\boldsymbol{\theta}_{t}),\nabla\mathcal{L}_{d}(\boldsymbol{\theta}_{t})\rangle<0$ then
11     $g_{t}^{dual}=\nabla\mathcal{(}\boldsymbol{\theta}_{t})_{\|(\nabla\mathcal{L}_{d}(\boldsymbol{\theta}_{t})^{\bot}}$
@@ 7     $g_{t}^{dual}=\nabla\mathcal{L}(\boldsymbol{\theta}_{t})$
In Algorithm 2, we first calculate a normalized sum of two directions $g^{c}_{t}$. $\angle\bigl(g^{c}_{t},\nabla\mathcal{L}_{d}(\boldsymbol{\theta}_{t})\bigr)=\angle\bigl(g^{c}_{t},\nabla\mathcal{L}_{p}(\boldsymbol{\theta}_{t})\bigr)=\frac{\phi_{t}}{2}$ 444we use $\angle\bigl(\mathbf{u},\mathbf{v}\bigr)$ to represent the angle between two vectors, mathematically: $\angle\bigl(\mathbf{u},\mathbf{v}\bigr):=\arccos\!\biggl(\frac{\mathbf{u}\cdot\mathbf{v}}{\|\mathbf{u}\|\;\|\mathbf{v}\|}\biggr)$ (16) , and then, we project the gradient of the total loss $\nabla\mathcal{L}(\theta_{t})$ onto $g^{c}_{t}$ to get the final common descent direction $g^{dual}_{t}$.
@@
$\nabla\mathcal{L}(\boldsymbol{\theta}_{t})=\nabla\mathcal{L}_{p}(\boldsymbol{\theta}_{t})+\nabla\mathcal{L}_{d}(\boldsymbol{\theta}_{t})$ (17)
@@
$\nabla\mathcal{L}(\boldsymbol{\theta}_{t})\in\mathbf{K}^{*}$ (18)
we use $\nabla\mathcal{L}(\boldsymbol{\theta}_{t})$ as the update. Otherwise, we first project $\nabla\mathcal{L}(\boldsymbol{\theta}_{t})$ onto two boundary $\nabla\mathcal{L}_{\|(\nabla\mathcal{L}_{p}(\boldsymbol{\theta}_{t}))^{\bot}}$ and $\nabla\mathcal{L}_{\|(\nabla\mathcal{L}_{d}(\theta_{t}))^{\bot}}$ of $\mathbf{G}_{t}$, respectively. Then, calculate the arithmetic average as the update:
$g^{dual}_{t}=\frac{1}{2}\nabla\mathcal{L}_{\|(\nabla\mathcal{L}_{p}(\boldsymbol{\theta}_{t}))^{\bot}}+\frac{1}{2}\nabla\mathcal{L}(\boldsymbol{\theta}_{t})_{\|(\nabla\mathcal{L}_{d}(\boldsymbol{\theta}_{t}))^{\bot}}$ (19)
This guarantees that the chosen descent direction reduces both $\mathcal{L}_{p}(\boldsymbol{\theta}_{t})$ and $\nabla\mathcal{L}_{d}(\boldsymbol{\theta}_{t})$ simultaneously.
In Algorithm 4, at iteration $t$, the dual gradient $g_{t}^{\mathrm{dual}}$ is constructed based on whether the current gradient of the total loss lies inside the dual cone $\mathbf{K}^{*}$ and whether it conflicts with bound gradients $\nabla\mathcal{L}_{d}(\boldsymbol{\theta}_{t})$ and $\nabla\mathcal{L}_{p}(\boldsymbol{\theta}_{t})$:
$g_{t}^{\mathrm{dual}}=\begin{cases}\nabla\mathcal{L}(\boldsymbol{\theta}_{t}),&\nabla\mathcal{L}(\boldsymbol{\theta}_{t})\in\mathbf{K}^{*},\\[4.30554pt] \displaystyle\nabla\mathcal{L}(\boldsymbol{\theta}_{t})_{\|(\nabla\mathcal{L}_{p}(\boldsymbol{\theta}_{t}))^{\bot}},&\nabla\mathcal{L}(\boldsymbol{\theta}_{t})\notin\mathbf{K}^{*},\ \langle\nabla\mathcal{L}(\boldsymbol{\theta}_{t}),\nabla\mathcal{L}_{p}(\boldsymbol{\theta}_{t})\rangle<0,\\[8.61108pt] \displaystyle\nabla\mathcal{(}\boldsymbol{\theta}_{t})_{\|(\nabla\mathcal{L}_{d}(\boldsymbol{\theta}_{t})^{\bot}},&\nabla\mathcal{L}(\boldsymbol{\theta}_{t})\notin\mathbf{K}^{*},\ \langle\nabla\mathcal{L}(\boldsymbol{\theta}_{t}),\nabla\mathcal{L}_{d}(\boldsymbol{\theta}_{t})\rangle<0.\end{cases}$ (20)
@@
Assume that both loss functions, $\mathcal{L}_{d}(\cdot)$ and $\mathcal{L}_{p}(\cdot)$, are differentiable and the total gradient $\nabla\mathcal{L}(\cdot)$ is $L$-Lipschitz continuous with $L>0$. If $g_{t}^{\text{dual}}$ satisfies the following two conditions:
@@
    $2\langle\nabla\mathcal{L}(\boldsymbol{\theta}_{t}),g_{t}^{\text{dual}}\rangle-\|g_{t}^{\text{dual}}\|^{2}\geq 0$,
@@
    There exists $M>0$ such that $\|g_{t}^{\text{dual}}\|\geq M\|\nabla\mathcal{L}(\boldsymbol{\theta}_{t})\|$,
@@
$\frac{1}{T+1}\sum_{t=0}^{T}\|\nabla\mathcal{L}(\boldsymbol{\theta}_{t})\|^{2}\leq\frac{2\left(\mathcal{L}(\boldsymbol{\theta}_{0})-\mathcal{L}(\boldsymbol{\theta}^{*})\right)}{\lambda M(T+1)}.$ (21)
@@ 3.2 Potential limitation of linear scalarization
In most contemporary PIML studies, the primary and most effective method for calculating the final total loss is through linear scalarization, as shown in the form illustrated in Eq 1. As is aware by many previous studies, the specific values of the coefficients $\alpha$ and $\beta$ cause significant difficulties in fine-tuning. Since the value of the coefficients will not change during the training process, it may lead to more serious problems:
@@
    Linear scalarization constructs a combined loss function as a linear weighted sum [^5] Here, we adopted a more general formulation, that is, we use $\theta$ to represent the entire parameter set and do not divide them into data-exclusive parameter sets and physics-exclusive parameter sets:
    $\mathcal{L}_{\alpha}(\boldsymbol{\theta})=\alpha\mathcal{L}_{1}(\boldsymbol{\theta})+(1-\alpha)\mathcal{L}_{2}(\boldsymbol{\theta}),\quad\alpha\in[0,1]$
@@
    $\nabla\mathcal{L}_{\alpha}(\boldsymbol{\theta}^{*})=(1-\alpha)\nabla\mathcal{L}_{2}(\boldsymbol{\theta}^{*})$
@@
    $\mathcal{L}_{\beta}(\theta)=\beta\mathcal{L}_{1}(\theta)+(1-\beta)\mathcal{L}_{2}(\theta)\quad\beta\in(0,1)$
@@
    However, certain Pareto stationary points found by MGDA cannot be theoretically achieved by optimizing the scalarized total loss for any interior linear scalarization weights $\beta\in(0,1)$.
@@
Figure 3: Difference between traditional training methods and training approaches based on multi-gradient descent
As illustrated in Figure 3, we provide a summary flow chart that highlights the fundamental difference between traditional training methods and training approaches based on multi-gradient descent.
On the left, the traditional method first constructs a linear scalarization loss function, typically of the form $\mathcal{L}(\boldsymbol{\theta})=\alpha\mathcal{L}_{\mathrm{data}}(\boldsymbol{\theta})+\beta\mathcal{L}_{\mathrm{physics}}(\boldsymbol{\theta})$, and then computes the gradient of this combined loss to update model parameters using standard gradient descent, where $\lambda$ is the step length, it could be either fixed or adaptive.
In contrast, the right side shows the pipeline of training approaches based on multi-gradient descent, which treats the data loss and physics loss as two distinct objectives. Rather than combining them into a single scalarized function, this method directly formulates a multi-objective optimization problem and computes a common descent direction $\mathbf{d}$ that ensures both losses decrease simultaneously. This approach eliminates the need for manual tuning of scalarization weights and enables more principled updates for all parameter types.
@@ 4.1 PIML model based on the macroscopic traffic flow model
The loss-function design for LWR-PINN is shown in Eq 25, where $N_{o}$ and $N_{a}$ denote the numbers of observation and auxiliary points, respectively.
For each observation point
@@
After tuning the hyperparameters, the experimental results are displayed in Table 5, where the LWR-PINN model refers to the model shown in Figure 4, which adopted a loss function as: $Loss_{\theta,\omega}^{LWR}=100\cdot MSE_{o}+500\cdot MSE_{a}$, for all other models, they treat the training process as a bi-objective optimization problem, the two objectives are $MSE_{o}$ and $MSE_{a}$. According to Table 5, the TMGD-LWR-PINN model and the models based on DCGDs exhibit performance similar to that of the baseline model, the LWR-PINN model, with the exception of the DCGD-AVER-LWR-PINN model, which shows a significantly larger speed $\mathbb{L}^{2}$ error compared to the baseline. In traffic flow theory Ni [2015], both traffic density $\rho(x,t)$ and speed $v(x,t)$ are functions of space $x$ and time $t$. In the context of macroscopic traffic flow models, field data typically refer to sensor measurements collected from detectors distributed along the highway. Mathematically, the sensor data consist of input-output pairs $\{(x_{i},t_{i}),(\rho_{i},v_{i})\}_{i=1}^{N}$, where each spatial-temporal coordinate $(x_{i},t_{i})$ corresponds to a unique measurement of density $\rho_{i}$ and speed $v_{i}$. In other words, the datasets $\mathbf{x}=\{(x_{i},t_{i})\}_{i=1}^{N}$ and $\mathbf{y}=\{(\rho_{i},v_{i})\}_{i=1}^{N}$ are aligned element-wise.
The PIML model first uses its data-driven component to approximate the density and speed functions. Based on these approximations, the physics residual is computed via automatic differentiation to enforce the macroscopic traffic flow model.
From the perspective of function approximation, high-resolution data are essential to accurately capture the underlying physical patterns. As illustrated in Figure 5, low-resolution data can lead to catastrophic errors in function approximation. However, most real-world traffic sensor data are inherently low-resolution, since the spatial resolution $\Delta x$ (i.e., the distance between two sensors) is often large, and the temporal resolution $\Delta t$ (i.e., the time interval between consecutive measurements) is usually no less than five minutes.
As a result, PIML may fail to approximate the true density function $\rho(x,t)$ due to insufficient data resolution Lei et al. [2025b]. In this case, the physics-informed term $\mathcal{L}_{\mathrm{physics}}$ becomes an invalid or misleading component in the overall loss function:
$\mathcal{L}=\alpha\mathcal{L}_{\mathrm{data}}+\beta\mathcal{L}_{\mathrm{physics}}$ (28)
whose gradient information is uninformative and can even misguide the optimization process.
Therefore, when the PIML model suffers from such an endogenous flaw induced by low-resolution training data, even an improved training strategy cannot yield better performance.
Figure 5: When low-resolution data meet the function approximation